You are in a 12 -member panel. If two people from the panel are selected to represent the panel, what is the probability you won't be selected?
Answer #1

Two selection means without replacement..,

Not selecting probability;

panel body 12.

1st time my choosen Probably is 1/12=0.084

and second time my chosen probably is 1/11=0.090

total probability by selecting me. Is 0.0274

but our que is not selecting me probability .

so

=1-0.0273

=0.9726